OpenStudy (darkbluechocobo):

The regular price of a pair of jeans is x dollars. Let f(x) = x − 14.60, where f represents the price of the jeans discounted by $14.60. Let g(x) = 0.9timesx represent the price of the jeans discounted by 10%. Use composition to find the price of the jeans if they are first discounted by $14.60, then by 10%.

OpenStudy (darkbluechocobo):

(g f)of(x) = 0.9timesx − 14.60 (f g)of(x) = 0.9timesx + 14.60 (f g)of(x) = 0.9timesx + 13.14 (g f)of(x) = 0.9timesx − 13.14

OpenStudy (darkbluechocobo):

@aum would be a?

OpenStudy (aum):

"first discounted by $14.60, then by 10%" implies calculate f(x) first and then g(x). So it is g(f(x)) or (g o f)(x). g(x) = 0.9x g(f(x)) = 0.9 * f(x) = 0.9(x - 14.60) = 0.9x - 13.14